# Stochastic Processes and Transport Plans

### v5

@@ 2. Concerning relation (63), we verify
First, we have $p_{t}\in C^{\infty}(\mathbb{R}^{d})$, $t\in[0,1)$ and since $p_{t}>0$ also $\log p_{t}\in C^{\infty}(\mathbb{R}^{d})$. Since $\nabla\|x\|^{2},\nabla f_{t}\in L^{2}(\mathbb{R}^{d},\mu_{t})$, we conclude that also $\nabla\log p_{t}\in L^{2}(\mathbb{R}^{d},\mu_{t})$. Furthermore, it is easy to see that $\lim_{r\to\infty}\max\{\log p_{t}(x):\|x\|\geq r\}\to-\infty$. Consequently, we obtain for functions $h_{n}\in C_{c}^{\infty}(\mathbb{R})$ that $\phi_{n}\coloneqq h_{n}\circ\log p_{t}\in C_{c}^{\infty}(\mathbb{R}^{d})$. Now consider especially functions with $h_{n}(s)=s$ for $-n\leq s\leq\|\log p_{t}\|_{\infty}$ and $|h_{n}^{\prime}|\leq 1$. Then we obtain with $C_{n}\coloneqq\{x\in\mathbb{R}^{d}:\log p_{t}(x)\geq-n\}$ that $\phi_{n}(x)=\log p_{t}(x)$ for $x\in C_{n}$ and thus
@@ 5 Curves via Markov kernels
Unfortunately, the approach of Lipman et al. [30], see Example 5.4, is not covered by the previous section. Indeed, we need a more general viewpoint via Markov kernels which is given in the following Subsection 5.1. We will see in Subsection 5.2 that curve-velocity pairs induced by couplings fit into this general setting.
@@ 5.1 General Construction
Assume that $\mathcal{K}_{t}(y,\cdot)$ is an absolutely continuous curve in $\mathcal{P}_{2}(\mathbb{R}^{d})$ for $\mu_{1}$ a.e. $y\in\mathbb{R}^{d}$ with a vector field $v_{t}^{y}$ such that ($\mathcal{K}_{t}(y,\cdot),v_{t}^{y})$ fulfills (CE) for a.e. $y$. Further, let $(t,x,y)\mapsto v_{t}^{y}$ be measurable and $\|v_{t}^{y}\|_{L^{2}(\mathbb{R}^{d}\times\mathbb{R}^{d},\alpha_{t})}\in L^{1}([0,1])$. Then $\alpha_{t}$ and $\nu_{t}$ are narrowly continuous. The curve $\nu_{t}$ is absolutely continuous and $(\nu_{t},v_{t})$ satisfies (CE), where
@@ 3. Finally, we conclude
Remark 5.3 (Relation to Lipman et al. [30]).
The authors in [30] argue only with densities. Then the notation of the so-called ”conditional probability path” translates as
@@
We recap the example in [30, Example II] with our approach.
@@
By [4, Lemma 7.1.10] we have that $W_{2}(\nu_{1},\mu_{1})\leq(1-r)\sqrt{\int\|x\|^{2}\mathrm{d}\mu_{1}}$ and thus $\nu_{1}$ converges to $\mu_{1}$ in $\mathcal{P}_{2}(\mathbb{R}^{d})$ as $r\to 1$. The authors of [30] then learn the velocity field $v_{t}$ of $\nu_{t}$ as explained in Example 7.2. $\diamond$
@@ 5.2 Curves induced by Couplings via Disintegration
Note that it is not a priori clear, that there are representatives of $v_{t}$ such that $v_{t}$ seen as a function $[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ is measurable. We show in Proposition B.1 why this is the case.
@@
i.e., $\nu_{t}=e_{t,\sharp}(\tilde{\alpha})$. By Example 7.2 the vector field induced by $\tilde{\alpha}$ coincides with the vector field constructed in (81).
@@ 6 Curves via Stochastic Processes
This section resembles the results from the papers of Liu et al. [31, 32] with our notation. As already mentioned in the introduction, these authors use a stochastic process $(X_{t})_{t}$ to determine an absolutely continuous curve $\mu_{t}$ and an associated vector field $v_{t}$ based on the conditional expectation
@@ 6.1 Conditional Expectation
Conditional expectations $\mathbb{E}[X|Y]$ of random vectors can also be interpreted as Borel measurable functions $\psi:\mathbb{R}^{d}\to\mathbb{R}^{d}$ via the Doob–Dynkin Lemma, see [25, Lemma 1.14].
@@ 6.2 Curve via Conditional Expectation
In [32, Theorem 3.3] it was proven that for well-behaved stochastic processes $(X_{t})_{t}$ the conditional expectation can be used to obtain a vector field $v_{t}$ such that $(X_{t,\sharp}\mathbb{P},v_{t})$ fulfills the continuity equation. For convenience, we include the proof.
@@
0 0 $=\int_{\Omega}(\varphi_{1}\circ X_{1}-\varphi_{0}\circ X_{0})\,\mathrm{d}\mathbb{P}=\int_{\Omega}\int_{0}^{1}\frac{\mathrm{d}}{\mathrm{d}t}(\varphi_{t}\circ X_{t})\,\mathrm{d}t\mathrm{d}\mathbb{P}$ (135)
@@
Let $X_{t}=(1-t)X_{0}+tX_{1}$ for independent random variables $X_{0},X_{1}$, where $X_{0}\sim\mathcal{N}(0,I_{d})$. Then in [47, Appendix A1] it is shown that the formula for the vector field associated with $\mu_{t}$,
@@
from Proposition 4.11 can also be derived via the velocity field from Theorem 6.3 and Tweedie’s formula. Recall that Tweedie’s formula [13, Lemma 3.2] for $X_{t}$ reads as
@@ 7 Flow Matching
The name ”FM” comes from ”flow matching”. In the original paper [30], the notation ”conditional flow matching” was used which refers to a ”conditional flow path”, see Remark 5.3. In practice, we work with the empirical expectation. To this end, we will need to be able to sample from $\alpha_{t}=\mathcal{K}_{t}(y,\cdot)\times_{y}\mu_{1}$ and to compute $v_{t}^{y}(x)$. A typical scenario is that samples $y_{i}$, $i=1,\ldots,n$ from the data distribution $\mu_{1}$ are available and that it is easy to sample from the Markov kernel $\mathcal{K}_{t}(y_{i},\cdot)$, $i=1,\ldots,n$.
@@
can be interpreted as a minimization problem. Namely, it is well-known, see, e.g., [8, 10.1.5 (5)], that
@@ 8.1 Flow Matching Algorithms
Given two measures $\mu_{0},\mu_{1}\in\mathcal{P}_{2}(\mathbb{R}^{d})$, sampling from an optimal transport plan $\alpha\in\Gamma_{o}(\mu_{0},\mu_{1})$ is usually not possible. Often both measures are not known analytically but only an approximation by samples, so that an optimal plan cannot be obtained. But even the plan for the approximation is often not computable since usually the number of samples is so high that solving the OT problem is computationally too expensive. In this case a popular method is minibatch OT flow matching [44]. The idea is to partition the samples $x_{1},\ldots,x_{N_{t}}\sim\mu_{1}$ into minibatches (smaller subsets), on which solving the OT problem is computational feasible. Details are described in Algorithm 2.
@@ 8.2 Numerical Examples
Figure 7: Trajectories of points from a vector field $u^{\theta}$ obtained via flow matching from a coupling $\alpha\in\Gamma(\mu_{0},\mu_{1})$. We chose $\mu_{0}=\mathcal{N}(0,1)$ and as target distribution $\mu_{1}$ a Gaussian mixture model with 888 modes. Black: points $\{z_{i}\}_{i=1}^{300}$ drawn from the source distribution $\mu_{0}=\mathcal{N}(0,1)$. Blue: points sampled via the vector field from $\{z_{i}\}_{i=1}^{300}$. Red: points sampled from the target distribution. Blue lines: trajectories of the vector field. The left image shows the results for $u^{\theta}$ trained using the independent coupling and the middle image shows the result when trained with minibatch OT flow matching. The right image shows the result from training a continuous normalizing flow as in Section 10.
Finally, we show a high dimensional application of flow matching in Figure 8. Here we used the flow vector field from [33], which was already trained via minibatch OT flow matching on images of cat faces. For a comparison of the performance for the flow matching with different plans see [44, Table 2].
Figure 8: Single trajectory for a flow matching model trained on cat images. For the trained model, see [33]. Note that for the standard Gaussian distribution we can sample in each direction separately, so that the left starting sample contains the one-dimensional coordinates in each of the $d=256^{2}$ directions for the red-green-blue channels.
@@ 9 Flow Matching in Bayesian Inverse Problems
In the next subsection, we will equip the space $\mathcal{P}_{P_{W}}$ with a metric turning it into a complete metric space and determine couplings such that the corresponding curve stays in $\mathcal{P}_{P_{W}}$. Then we will use these curves for flow matching in order to learn conditional generating networks that approximate the posterior distribution $P_{X|W=w}$ for $w$ sampled from $P_{W}$. This section is based on our paper [9], see also [26].
@@ 9.1 Conditional Wasserstein Distance
Conditional Wasserstein distances appear already in [28, Section 2.4], [34] as well as in the PhD thesis of Gigli [15, Chapter 4], who studied a metric on the so-called ”geometric tangent space” of the Wasserstein-222 space. Recently, conditional Wasserstein distances were studied in the realm of machine learning in [6] as well as in functional optimal transport in [23]. We will focus on its application in Bayesian flow matching as presented in our paper [9], where also the proofs of the following theorems and propositions are provided.
@@
Now we define the conditional Wasserstein(-222) distance by
@@
[…]
As shown in [26, Theorem 3,4] and [6], there is a counterpart of Theorem 3.3 for the conditional Wasserstein distance.
@@
To proof $\rm{i)}$ and $\rm{ii)}$ follow from [9, Lemma 5, Proposition 6]. The first statements of $\rm{iii)}$ follow from Proposition 4.6 and finally the absolutely continuity with respect to $W_{2,\eta}$ is implied by Theorem 9.3. ∎
@@
The counterpart of Proposition 4.4 reads as follows, see [9, Lemma 5]).
@@ 9.2 Almost Conditional Couplings
One drawback of the space $\mathcal{P}_{\eta}(\mathbb{R}^{m}\times\mathbb{R}^{d})$ is that we can in general not approximate $\mu\in\mathcal{P}_{\eta}(\mathbb{R}^{m}\times\mathbb{R}^{d})$ by an empirical measure, if $\eta$ is not empirical. In other words, an empirical approximation $\mu_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{(w_{i},x_{i})}$ of $\mu\in\mathcal{P}_{\eta}(\mathbb{R}^{m}\times\mathbb{R}^{d})$ with respect to $W_{2}$ is usually not an element of $\mathcal{P}_{\eta}(\mathbb{R}^{m}\times\mathbb{R}^{d})$. Even if we approximate two measures $\mu,\nu\in\mathcal{P}_{\eta}(\mathbb{R}^{m}\times\mathbb{R}^{d})$ with respect to $W_{2}$ by empirical measures $\mu_{n},\nu_{n}$ with $\pi^{1}_{\sharp}\mu_{n}=\pi^{1}_{\sharp}\nu_{n}$, it is not necessarily true that there is a sequence $\alpha_{n}\in\Gamma_{\pi^{1}_{\sharp}\mu_{n},o}(\mu_{n},\nu_{n})$ such that $\alpha_{n}$ converges to an $\alpha\in\Gamma_{\eta,o}(\mu,\nu)$ as [9, Example 9] shows.
@@
defines a metric on $\mathcal{P}_{2}(\mathbb{R}^{m}\times\mathbb{R}^{d})$. It turns out that for $0\ll\beta$ the optimal couplings $\alpha$ in this metric are ”close” to fulfilling $\pi^{1,1}\alpha=\Delta\eta$ in the following sense, see [9, Proposition 10].
@@
The distance $W_{\beta}$ successfully addresses the issue of approximating measures by empirical measures as the following proposition from [9, Proposition 12] shows.

### v6

@@ 2. Concerning relation (63), we verify
First, we have $p_{t}\in C^{\infty}(\mathbb{R}^{d})$, $t\in[0,1)$ and since $p_{t}>0$ also $\log p_{t}\in C^{\infty}(\mathbb{R}^{d})$. Since $\nabla\|x\|^{2},\nabla f_{t}\in L^{2}(\mathbb{R}^{d},\mu_{t})$, we conclude that also $\nabla\log p_{t}\in L^{2}(\mathbb{R}^{d},\mu_{t})$. Furthermore, it is easy to see that $\lim_{r\to\infty}\max\{\log p_{t}(x):\|x\|\geq r\}\to$-$\infty$. Consequently, we obtain for functions $h_{n}\in C_{c}^{\infty}(\mathbb{R})$ that $\phi_{n}\coloneqq h_{n}\circ\log p_{t}\in C_{c}^{\infty}(\mathbb{R}^{d})$. Now consider especially functions with $h_{n}(s)=s$ for $n\leq s\leq\|\log p_{t}\|_{\infty}$ and $|h_{n}^{\prime}|\leq 1$. Then we obtain with $C_{n}\coloneqq\{x\in\mathbb{R}^{d}:\log p_{t}(x)\geq-n\}$ that $\phi_{n}(x)=\log p_{t}(x)$ for $x\in C_{n}$ and thus
@@ 5 Curves via Markov kernels
Unfortunately, the approach of Lipman et al. [lipman2023flow], see Example 5.4, is not covered by the previous section. Indeed, we need a more general viewpoint via Markov kernels which is given in the following Subsection 5.1. We will see in Subsection 5.2 that curve-velocity pairs induced by couplings fit into this general setting.
@@ 5.1 General Construction
Assume that $\mathcal{K}_{t}(y,\cdot)$ is an absolutely continuous curve in $\mathcal{P}_{2}(\mathbb{R}^{d})$ for $\mu_{1}$ - a.e. $y\in\mathbb{R}^{d}$ with a vector field $v_{t}^{y}$ such that ($\mathcal{K}_{t}(y,\cdot),v_{t}^{y})$ fulfills (CE) for a.e. $y$. Further, let $(t,x,y)\mapsto v_{t}^{y}$ be measurable and $\|v_{t}^{y}\|_{L^{2}(\mathbb{R}^{d}\times\mathbb{R}^{d},\alpha_{t})}\in L^{1}([0,1])$. Then $\alpha_{t}$ and $\nu_{t}$ are narrowly continuous. The curve $\nu_{t}$ is absolutely continuous and $(\nu_{t},v_{t})$ satisfies (CE), where
@@ 3. Finally, we conclude
Remark 5.3 (Relation to Lipman et al. [lipman2023flow]).
The authors in [lipman2023flow] argue only with densities. Then the notation of the so-called ”conditional probability path” translates as
@@
We recap the example in [lipman2023flow, Example II] with our approach.
@@
By [AGS2008, Lemma 7.1.10] we have that $W_{2}(\nu_{1},\mu_{1})\leq(1-r)\sqrt{\int\|x\|^{2}\mathrm{d}\mu_{1}}$ and thus $\nu_{1}$ converges to $\mu_{1}$ in $\mathcal{P}_{2}(\mathbb{R}^{d})$ as $r\to 1$. The authors of [lipman2023flow] then learn the velocity field $v_{t}$ of $\nu_{t}$ as explained in Example 7.2. $\diamond$
@@ 5.2 Curves induced by Couplings via Disintegration
Note that it is not a priori clear, that there are representatives of $v_{t}$ such that $v_{t}$ seen as a function $0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ is measurable. We show in Proposition B.1 why this is the case.
@@
[…]
@@ 6 Curves via Stochastic Processes
This section resembles the results from the papers of Liu et al. [liu2022rectified, liu2023flow] with our notation. As already mentioned in the introduction, these authors use a stochastic process $(X_{t})_{t}$ to determine an absolutely continuous curve $\mu_{t}$ and an associated vector field $v_{t}$ based on the conditional expectation
@@ 6.1 Conditional Expectation
Conditional expectations $\mathbb{E}[X|Y]$ of random vectors can also be interpreted as Borel measurable functions $\psi:\mathbb{R}^{d}\to\mathbb{R}^{d}$ via the Doob–Dynkin Lemma, see [kallenberg1997foundations, Lemma 1.14].
@@ 6.2 Curve via Conditional Expectation
In [liu2023flow, Theorem 3.3] it was proven that for well-behaved stochastic processes $(X_{t})_{t}$ the conditional expectation can be used to obtain a vector field $v_{t}$ such that $(X_{t,\sharp}\mathbb{P},v_{t})$ fulfills the continuity equation. For convenience, we include the proof.
@@
$0$ $=\int_{\Omega}(\varphi_{1}\circ X_{1}-\varphi_{0}\circ X_{0})\,\mathrm{d}\mathbb{P}=\int_{\Omega}\int_{0}^{1}\frac{\mathrm{d}}{\mathrm{d}t}(\varphi_{t}\circ X_{t})\,\mathrm{d}t\mathrm{d}\mathbb{P}$ (135)
@@
Let $X_{t}=(1-t)X_{0}+tX_{1}$ for independent random variables $X_{0},X_{1}$, where $X_{0}\sim\mathcal{N}(0,I_{d})$. Then in [zhang2024flow, Appendix A1] it is shown that the formula for the vector field associated with $\mu_{t}$,
@@
from Proposition 4.11 can also be derived via the velocity field from Theorem 6.3 and Tweedie’s formula. Recall that Tweedie’s formula [daras2024consistent, Lemma 3.2] for $X_{t}$ reads as
@@ 7 Flow Matching
The name ”FM” comes from ”flow matching”. In the original paper [lipman2023flow], the notation ”conditional flow matching” was used which refers to a ”conditional flow path”, see Remark 5.3. In practice, we work with the empirical expectation. To this end, we will need to be able to sample from $\alpha_{t}=\mathcal{K}_{t}(y,\cdot)\times_{y}\mu_{1}$ and to compute $v_{t}^{y}(x)$. A typical scenario is that samples $y_{i}$, $i=1,\ldots,n$ from the data distribution $\mu_{1}$ are available and that it is easy to sample from the Markov kernel $\mathcal{K}_{t}(y_{i},\cdot)$, $i=1,\ldots,n$.
@@
can be interpreted as a minimization problem. Namely, it is well-known, see, e.g., [bogachev2007measure, 10.1.5 (5)], that
@@ 8.1 Flow Matching Algorithms
Given two measures $\mu_{0},\mu_{1}\in\mathcal{P}_{2}(\mathbb{R}^{d})$, sampling from an optimal transport plan $\alpha\in\Gamma_{o}(\mu_{0},\mu_{1})$ is usually not possible. Often both measures are not known analytically but only an approximation by samples, so that an optimal plan cannot be obtained. But even the plan for the approximation is often not computable since usually the number of samples is so high that solving the OT problem is computationally too expensive. In this case a popular method is minibatch OT flow matching [tong2023improving]. The idea is to partition the samples $x_{1},\ldots,x_{N_{t}}\sim\mu_{1}$ into minibatches (smaller subsets), on which solving the OT problem is computational feasible. Details are described in Algorithm 2.
@@ 8.2 Numerical Examples
Figure 7: Trajectories of points from a vector field $u^{\theta}$ obtained via flow matching from a coupling $\alpha\in\Gamma(\mu_{0},\mu_{1})$. We chose $\mu_{0}=\mathcal{N}(0,1)$ and as target distribution $\mu_{1}$ a Gaussian mixture model with 88 modes. Black: points $\{z_{i}\}_{i=1}^{300}$ drawn from the source distribution $\mu_{0}=\mathcal{N}(0,1)$. Blue: points sampled via the vector field from $\{z_{i}\}_{i=1}^{300}$. Red: points sampled from the target distribution. Blue lines: trajectories of the vector field. The left image shows the results for $u^{\theta}$ trained using the independent coupling and the middle image shows the result when trained with minibatch OT flow matching. The right image shows the result from training a continuous normalizing flow as in Section 10.
Finally, we show a high dimensional application of flow matching in Figure 8. Here we used the flow vector field from [martin2024pnp], which was already trained via minibatch OT flow matching on images of cat faces. For a comparison of the performance for the flow matching with different plans see [tong2023improving, Table 2].
Figure 8: Single trajectory for a flow matching model trained on cat images. For the trained model, see [martin2024pnp]. Note that for the standard Gaussian distribution we can sample in each direction separately, so that the left starting sample contains the one-dimensional coordinates in each of the $d=256^{2}$ directions for the red-green-blue channels.
@@ 9 Flow Matching in Bayesian Inverse Problems
In the next subsection, we will equip the space $\mathcal{P}_{P_{W}}$ with a metric turning it into a complete metric space and determine couplings such that the corresponding curve stays in $\mathcal{P}_{P_{W}}$. Then we will use these curves for flow matching in order to learn conditional generating networks that approximate the posterior distribution $P_{X|W=w}$ for $w$ sampled from $P_{W}$. This section is based on our paper [chemseddine2024conditional], see also [kerrigan2024dynamic].
@@ 9.1 Conditional Wasserstein Distance
Conditional Wasserstein distances appear already in [kloeckner2021extensions, Section 2.4], [peszek2023heterogeneous] as well as in the PhD thesis of Gigli [gig, Chapter 4], who studied a metric on the so-called ”geometric tangent space” of the Wasserstein-$2$ space. Recently, conditional Wasserstein distances were studied in the realm of machine learning in [barboni2024understanding] as well as in functional optimal transport in [hosseini2024conditional]. We will focus on its application in Bayesian flow matching as presented in our paper [chemseddine2024conditional], where also the proofs of the following theorems and propositions are provided.
@@
Now we define the conditional Wasserstein(-$2$) distance by
@@
Proposition 9.1ii)ii) indicates another use of the conditional Wasserstein distance. Namely if we have two measures $\mu_{0},\mu_{1}\in\mathcal{P}_{\eta}$ and we have that $W_{2}(\mu_{0},\mu_{1})$ is small that does not necessarily mean that the posteriors are close. We can see in Figure 9 that there exist $\eta\in\mathcal{P}_{2}(\mathbb{R}),\mu_{n},\mu_{n}\in\mathcal{P}_{\eta}(\mathbb{R}\times\mathbb{R})$ such that $W_{2}(\mu_{n},\nu_{n})=1$ and $\mathbb{E}_{w\sim\eta}[W_{2}^{2}(\mu_{0,w},\mu_{1,w})]=W_{2,\eta}(\mu_{n},\nu_{n})=n$. This means that $W_{2}$ is not a good measure in order to access if a generator $g$ is approximating the posteriors well. In contrast if we use $W_{2,\eta}$ we know that the posteriors are, at least in expectation, well approximated.
As shown in [kerrigan2024dynamic, Theorem 3,4] and [barboni2024understanding], there is a counterpart of Theorem 3.3 for the conditional Wasserstein distance.
@@
To proof $\rm{i)}$ and $\rm{ii)}$ follow from [chemseddine2024conditional, Lemma 5, Proposition 6]. The first statements of $\rm{iii)}$ follow from Proposition 4.6 and finally the absolutely continuity with respect to $W_{2,\eta}$ is implied by Theorem 9.3. ∎
@@
The counterpart of Proposition 4.4 reads as follows, see [chemseddine2024conditional, Lemma 5]).
@@ 9.2 Almost Conditional Couplings
One drawback of the space $\mathcal{P}_{\eta}(\mathbb{R}^{m}\times\mathbb{R}^{d})$ is that we can in general not approximate $\mu\in\mathcal{P}_{\eta}(\mathbb{R}^{m}\times\mathbb{R}^{d})$ by an empirical measure, if $\eta$ is not empirical. In other words, an empirical approximation $\mu_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{(w_{i},x_{i})}$ of $\mu\in\mathcal{P}_{\eta}(\mathbb{R}^{m}\times\mathbb{R}^{d})$ with respect to $W_{2}$ is usually not an element of $\mathcal{P}_{\eta}(\mathbb{R}^{m}\times\mathbb{R}^{d})$. Even if we approximate two measures $\mu,\nu\in\mathcal{P}_{\eta}(\mathbb{R}^{m}\times\mathbb{R}^{d})$ with respect to $W_{2}$ by empirical measures $\mu_{n},\nu_{n}$ with $\pi^{1}_{\sharp}\mu_{n}=\pi^{1}_{\sharp}\nu_{n}$, it is not necessarily true that there is a sequence $\alpha_{n}\in\Gamma_{\pi^{1}_{\sharp}\mu_{n},o}(\mu_{n},\nu_{n})$ such that $\alpha_{n}$ converges to an $\alpha\in\Gamma_{\eta,o}(\mu,\nu)$ as [chemseddine2024conditional, Example 9] shows.
@@
defines a metric on $\mathcal{P}_{2}(\mathbb{R}^{m}\times\mathbb{R}^{d})$. It turns out that for $0\ll\beta$ the optimal couplings $\alpha$ in this metric are ”close” to fulfilling $\pi^{1,1}\alpha=\Delta\eta$ in the following sense, see [chemseddine2024conditional, Proposition 10].
@@
The distance $W_{\beta}$ successfully addresses the issue of approximating measures by empirical measures as the following proposition from [chemseddine2024conditional, Proposition 12] shows.